OpenStudy (yoloshroom):

determine inverse function of g(t)=7^(t-2)

OpenStudy (yoloshroom):

first we do t=y^(y-2) but then what? explain whole process plllss

OpenStudy (pawanyadav):

Y=7^(t-2) Convert t as a function of y. Then replace t and y with each other The new function will be inverse .

OpenStudy (yoloshroom):

SO t=7^(Y-2) I i need to solve for y, how can i do that?

OpenStudy (pawanyadav):

Take log both side logy=(t-2)log7 logy/log7=t-2 So t=2+logy/log7 Now replace t by y and y by t...you will gget y=logt/log7+2 Is inverse of that.

OpenStudy (yoloshroom):

Ah, thank you.